# Joint Beamforming and Antenna Position Optimization for Fluid Antenna-Assisted MU-MIMO Networks

### v1

@@ Abstract
The fluid antenna system (FAS) has emerged as a disruptive technology for future wireless networks, offering unprecedented degrees of freedom (DoF) through the dynamic configuration of antennas in response to propagation environment variations. The integration of fluid antennas (FAs) with multiuser multiple-input multiple-output (MU-MIMO) networks promises substantial weighted sum rate (WSR) gains via joint beamforming and FA position optimization. However, the joint design is challenging due to the strong coupling between beamforming matrices and antenna positions. To address the challenge, we propose a novel block coordinate ascent (BCA)-based method in FA-assisted MU-MIMO networks. Specifically, we first employ matrix fractional programming techniques to reformulate the original complex problem into a more tractable form. Then, we solve the reformulated problem following the BCA principle, where we develop a low-complexity majorization maximization algorithm capable of optimizing all FA positions simultaneously. To further reduce the computational, storage, and interconnection costs, we propose a decentralized implementation for our proposed algorithm by utilizing the decentralized baseband processing (DBP) architecture. Simulation results demonstrate that with our proposed algorithm, the FA-assisted MU-MIMO system achieves up to a $47\%$ WSR improvement over conventional MIMO networks equipped with fixed-position antennas. Moreover, the decentralized implementation reduces computation time by approximately $70\%$ and has similar performance compared with the centralized implementation.
Index Terms:
Fluid antenna system (FAS), MU-MIMO, fractional programming (FP), majorization maximization (MM), decentralized baseband processing (DBP).
@@ I Introduction
The sixth-generation (6G) wireless communication systems are expected to revolutionize global connectivity by enabling applications such as digital twins, precision healthcare, edge intelligence, and extended reality (XR) [1, 2]. These applications impose unprecedented requirements including terabit-per-second data rates, high energy efficiency, and sub-millisecond latency [3, 4]. To achieve these objectives, massive multiple-input multiple-output (MIMO) and multiuser MIMO (MU-MIMO) technology [5, 6] will play pivotal roles. The major advantage of MIMO systems is their ability to leverage spatial degrees of freedom (DoF), which can improve the system performance to exploit spatial diversity and multiplexing gains [7]. However, conventional MIMO systems typically assume fixed-position antenna (FPA) configurations, which lack the flexibility to dynamically adapt to changing propagation environments and thus cannot fully exploit the spatial DoF. Specifically, some antennas inevitably experience deep fading and may suffer from high interference, leading to significant performance loss.
To address these challenges, the fluid antenna system (FAS) was proposed in [8] as a disruptive technology. FAS utilizes fluid antennas (FAs) to achieve flexible control over the positions, gains, radiation patterns, and other characteristics of the antennas [9]. Among these techniques, position optimization has been an efficient way to fully harness the spatial DoF. The advantages of utilizing FAs in MU-MIMO networks are twofold. On one hand, the position of the FAs can be changed to avoid deep fading channels for the desirable links, thereby improving the optimal diversity-multiplexing tradeoff [10]. On the other hand, FAs can mitigate multiuser interference (MUI) by optimizing antenna positions because interference experiences deep fade with proper adjustments [11, 12, 13].
I-A Related Work
The concept of FAS was first introduced to wireless communications in [8]. Shortly after, a more precise definition of FAS and a channel model for single-antenna discrete port selection were proposed in [9]. The FAS with discrete port selection can be implemented by switchable pixels [14], and demonstrated superior performance over FPA systems [15, 16, 17, 18, 19, 10]. Although the discrete design simplifies hardware implementation, its performance is limited by the granularity of predefined ports. To address the challenge, an FAS with continuous position optimization implemented by liquid materials was proposed [20, 21]. However, such FAS still fails to fully harness spatial DoF, as it only allows a single FA to move along a one-dimensional line. Thus, the implementation of motor-driven FAS was later introduced in [22, 23], enabling continuous position optimization and seamless integration into MIMO systems. The advantages of motor-driven FAS were demonstrated in [24], where the authors analyzed an FA-assisted point-to-point MIMO system. By deriving the outage probability in closed form and numerically computing the system capacity gain, the authors proved that the motor-driven FAS offers higher stability and an increased data rate. Therefore, we focus on the motor-driven FAS in this paper.
Optimizing FA positions is key to unlocking the full potential of FAS. The first FA position optimization algorithm was introduced in [23]. The authors addressed the FA position optimization in a point-to-point MIMO system by formulating a non-concave optimization problem and solving it with a successive concave approximation (SCA)-based algorithm. However, this work did not consider joint beamforming and FA position optimization. Due to the strong coupling between beamforming matrices and FA positions, the SCA-based approach cannot be directly applied to this joint optimization problem. To address this challenge, the authors of [25] studied a secure rate maximization problem with joint beamforming, FA position control, and artificial noise optimization problem. Then, they proposed the minimum mean square error (MMSE) method to decouple the optimization variables, demonstrating that the joint optimization significantly improves performance. The FAS was further extended to multiuser systems by [12], where the authors considered the transmit power minimization problem in an FA-assisted multiuser system with joint beamforming and FA position optimization. They decoupled the variables using zero-forcing (ZF) and MMSE methods. Then, they optimized FA positions via a multi-directional descent (MDD) framework. It was demonstrated that FAS can effectively cancel MUI, thereby significantly reducing transmit power while ensuring the data rate. More recently, [26] evaluated the system capacity gain of the FAS by solving the weighted sum rate (WSR) maximization problem in an FA-assisted downlink MU-multiple-input single-output (MISO) system. The authors first applied the weighted minimum mean square error (WMMSE) algorithm to decouple beamforming vectors and FA positions. Then, they leveraged the majorization maximization (MM) framework to solve the resulting non-concave FA position optimization problem. Simulation results demonstrated that FAS enhances overall system capacity in MU-MISO networks.
However, the MM framework in [26] has unaffordable complexity when the number of FAs at the BS is large. Specifically, the MM framework optimizes one FA position at a time while keeping the others fixed in each iteration, leading to the increased computational time as the number of FAs grows.
I-B Contributions
@@
In this paper, we propose a novel block coordinate ascent (BCA)-based algorithm for joint beamforming and FA position optimization in the FA-assisted downlink MU-MIMO system, as well as provide a decentralized implementation of the proposed algorithm. Our contributions are summarized as follows.
@@
  • Different from [26] that focuses on MU-MISO networks, we consider the more general FA-assisted MU-MIMO networks. We formulate the joint beamforming and antenna position optimization as a WSR maximization problem, where the objective function is non-concave and the optimization variables are highly coupled. To decouple the beamforming matrices and FA positions, we utilize two matrix fractional programming (FP) techniques, i.e., the quadratic transform and the Lagrangian dual transform [27, 28]. These FP techniques enable us to solve the decoupled subproblems by a BCA-based algorithm.
  • Unlike existing FA position optimization algorithms that sequentially update FA positions, we propose a novel optimization algorithm based on the MM framework that optimizes all FA positions simultaneously. The proposed algorithm reduces computational time by exploiting parallelism and can be easily extended to a decentralized implementation. Additionally, we introduce a box-constrained movement mode to simplify both the formulated problem and the engineering implementation of FAS. This movement mode provides closed-form solutions to reduce complexity and enables decentralized implementation by allowing the independent movement of each FA.
@@
  • To further reduce computational, storage, and interconnection costs, we propose a decentralized implementation of our algorithm by utilizing the decentralized baseband processing (DBP) architecture [29], which partitions the transmit FA array into several clusters. The DBP architecture can decompose the optimization problem into smaller subproblems, and enable decentralized units (DUs) to solve them in parallel. To enable the decentralized implementation of the proposed BCA-based algorithm, we employ the non-homogeneous transform and Nesterov’s extrapolation [30, 31] to avoid the matrix inversion in the optimization of beamforming matrices. The decentralized implementation can significantly alleviate computational, storage, and interconnection costs with negligible performance loss.
  • We conduct various numerical experiments to verify the performance of our proposed algorithm. It is shown that the proposed algorithm can improve the WSR of FA-assisted MU-MIMO networks by $47\%$ compared with conventional MIMO networks using FPAs. Moreover, the decentralized implementation reduces computation time by approximately $70\%$ and has similar performance compared with the centralized implementation.
@@
The remainder of this paper is organized as follows. Section II presents the channel model of the FA-assisted MU-MIMO system and formulates the WSR problem. Section III reformulates the problem using FP techniques and solves it using BCA and MM. The decentralized implementation of the proposed algorithm is introduced in Section IV. Simulation results are provided in Section V, and conclusions are drawn in Section VI.
[…]
II System Model and Problem Formulation
As shown in Fig. 1, we consider a downlink MU-MIMO system where a BS with $M$ FAs serves $K$ users, each equipped with $N$ FAs. A three-dimensional (3D) Cartesian coordinate system is established to describe the positions of the transmit FAs at the BS and receive FAs at the users. Specifically, let $\mathbf{t}_{m}=\left[x_{m}^{t},y_{m}^{t},z_{m}^{t}\right]^{\mathsf{T}}\in\mathcal{C}_{m}^{t}$, $1\leq m\leq M$, denote the position of $m$-th transmit FA at the BS and let $\mathbf{r}_{kn}=\left[x_{kn}^{r},y_{kn}^{r},z_{kn}^{r}\right]^{\mathsf{T}}\in\mathcal{C}_{kn}^{r}$, $1\leq k\leq K$, $1\leq n\leq N$, denote the position of the $n$-th receive FA at user $k$, where $\mathcal{C}_{m}^{t}$ and $\mathcal{C}_{kn}^{r}$ are the given 3D movable regions of transmit and receive FAs. Without loss of generality, the movable regions are assumed to be cuboid [12], i.e., $\mathcal{C}_{m}^{t}=\left[x_{m}^{\min},x_{m}^{\max}\right]\times\left[y_{m}^{\min},y_{m}^{\max}\right]\times\left[z_{m}^{\min},z_{m}^{\max}\right]$ and $\mathcal{C}_{kn}^{r}=\left[x_{kn}^{\min},x_{kn}^{\max}\right]\times\left[y_{kn}^{\min},y_{kn}^{\max}\right]\times\left[z_{kn}^{\min},z_{kn}^{\max}\right]$, $\forall k,n$, where we assume the FA movable regions of the $K$ users to be identical.
@@ II System Model and Problem Formulation
In this paper, we consider narrow-band slow fading channels, where the transmit and receive FAs remain static or move slowly within the movable region during each quasi-static fading block. Moreover, the time required for FA movement is assumed to be much shorter than the coherence time [23]. Let $\mathbf{s}_{k}\in\mathbb{C}^{d}$ denote the data stream desired by user $k$, where $d\leq\min\{M,N\}$ is the number of parallel data streams, and assume that $\mathbf{s}_{k}\sim\mathcal{CN}\left(\mathbf{0},\mathbf{I}\right)$. Define $\mathbf{W}_{k}\in\mathbb{C}^{M\times d}$ as the beamforming matrix for transmitting data $\mathbf{s}_{k}$ from the BS to user $k$. The received signal $\mathbf{y}_{k}$ at user $k$ is given by

### v2

@@ Abstract
The fluid antenna system (FAS) is a disruptive technology for future wireless communication networks. This paper considers the joint optimization of beamforming matrices and antenna positions for weighted sum rate (WSR) maximization in fluid antenna (FA)-assisted multiuser multiple-input multiple-output (MU-MIMO) networks, which presents significant challenges due to the strong coupling between beamforming and FA positions, the non-concavity of the WSR objective function, and high computational complexity. To address these challenges, we first propose a novel block coordinate ascent (BCA)-based method that employs matrix fractional programming techniques to reformulate the original complex problem into a more tractable form. Then, we develop a parallel majorization maximization (MM) algorithm capable of optimizing all FA positions simultaneously. To further reduce computational costs, we propose a decentralized implementation based on the decentralized baseband processing (DBP) architecture. Simulation results demonstrate that our proposed algorithm not only achieves significant WSR improvements over conventional MIMO networks but also outperforms the existing method. Moreover, the decentralized implementation substantially reduces computation time while maintaining similar performance compared with the centralized implementation.
@@ I Introduction
The sixth-generation (6G) wireless communication systems aim to achieve terabit-per-second data rates, high energy efficiency, and sub-millisecond latency [2, 3, 4]. To achieve these objectives, massive multiple-input multiple-output (MIMO) and multiuser MIMO (MU-MIMO) technologies [5, 6] will play pivotal roles. The major advantage of MIMO systems is their ability to leverage spatial degrees of freedom (DoF), which can improve the system performance by exploiting spatial diversity and multiplexing gains [7]. However, conventional MIMO systems typically assume fixed-position antenna (FPA) configurations, which lack adaptability to varying propagation environments and thus cannot fully exploit the spatial DoF. Specifically, some antennas inevitably suffer from deep fading and may be subject to strong interference, leading to significant degradation in the signal-to-interference-plus-noise ratio (SINR).
To address these challenges, the fluid antenna system (FAS) was proposed in [8] as a disruptive technology. FASs leverage fluid antennas (FAs) to enable flexible control over antenna positions, gains, radiation patterns, and other key characteristics [9]. Among these capabilities, position reconfiguration has proven to be an effective means of fully exploiting spatial DoF. Existing implementations of position-reconfigurable FASs include pixel-based [10, 11, 12], liquid-based [13, 14], and motor-driven designs [15, 16, 17].
Position-reconfigurable FASs can be broadly categorized into pixel-based [10, 11, 12], liquid-based [13, 14], and motor-driven implementations [15, 16, 17]. Among these categories, motor-driven FASs offer higher reconfiguration fidelity and can be seamlessly integrated into MIMO systems. Considering the movement delay of motor-driven FASs, it is primarily envisioned to be deployed in ultra massive machine-type communication (umMTC) scenarios where the surrounding environment varies slowly [18]. [^1] Motor-driven FASs also face other implementation issues, including potentially excessive motor power consumption and calibration errors. Nevertheless, recent works have attempted to alleviate movement delay [19] and to jointly optimize motor power consumption [20]. We do not consider these implementation issues, as the focus of this paper is on analyzing the performance limits of FASs. The advantages of incorporating FAs into MU-MIMO networks are twofold. First, the positions of FAs can be dynamically adjusted to avoid deep fading in desired links, thereby enhancing the optimal diversity–multiplexing tradeoff [21]. Second, FAs help mitigate multiuser interference (MUI) by optimizing antenna locations, since interference can experience deep fading with appropriate position adjustments [22, 23, 24].
I-A Motivation
The weighted sum rate (WSR) maximization problem is essential to optimize the overall system capacity in MU-MIMO systems [25, 26], which has been extensively studied in the context of conventional MIMO systems [27, 28, 29]. However, the WSR maximization problem involving joint beamforming and FA position optimization in FA-assisted MU-MIMO systems poses significant and unprecedented challenges. First, beamforming and FA positions are strongly coupled. On one hand, beamforming alters the effective path coefficients of the channels, making the optimal FA positions dependent on the beamforming matrices. On the other hand, the array manifold, which is directly affected by the FA positions, plays a critical role in shaping the array beam pattern [30]. Although the method in [31] effectively decouples beamforming and FA positions, it cannot be readily extended to MU-MIMO setups. Second, the relationship between the channel characteristics and FA positions is highly non-linear, rendering the WSR maximization problem non-convex. In particular, small-scale fading induces rapid and irregular variations in the channel with respect to (w.r.t.) the FA positions [17], posing significant challenges for accurately determining the optimal FA configuration. To cope with such non-convexity, existing works often resort to iterative optimization techniques such as successive concave approximation (SCA) [16] and majorization–maximization (MM) [32, 31]. However, their performance is largely limited by the tightness of the designed surrogate functions. Last but not least, as a consequence of the non-convexity of the optimization problem, the resulting algorithms are computationally intensive, which hinders scalability to massive MIMO scenarios. These significant challenges highlight the need for novel approaches to maximize the WSR in FA-assisted MU-MIMO systems, thereby motivating the joint design of beamforming and antenna positions investigated in this work.
I-B Contribution
@@
In this paper, we propose a novel algorithm for joint beamforming and FA position optimization in FA-assisted downlink MU-MIMO systems within the block coordinate ascent (BCA) framework. In addition, a decentralized implementation is developed to further reduce computational overhead. The main contributions are summarized as follows.
@@
  • In contrast to [31], which focused on MU-multiple-input single-output (MU-MISO) networks, this work considers the more general FA-assisted MU-MIMO networks. We formulate the joint beamforming and antenna position optimization as a WSR maximization problem, where the objective function is non-concave and the optimization variables are highly coupled. To decouple the beamforming matrices and FA positions, we utilize two matrix fractional programming (FP) techniques, i.e., the quadratic transform and the Lagrangian dual transform [27, 28]. [^2] It has been shown in [27, 28] that the WMMSE approach used in [31] is equivalent to the FP techniques adopted in this paper. These FP techniques enable a BCA-based algorithm to solve the decoupled subproblems efficiently.
  • Unlike existing FA position optimization algorithms that sequentially update FA positions, we propose a novel parallel optimization algorithm based on the MM framework that simultaneously optimizes all FA positions. A tight surrogate function is constructed using matrix chain rules, providing a more accurate approximation of the original objective. The proposed algorithm outperforms the method in [31] and can be readily extended to a decentralized implementation to further reduce computational complexity.
@@
  • To further reduce computational cost, we propose a decentralized implementation of the algorithm using the decentralized baseband processing (DBP) architecture [33], which partitions the transmit FA array into multiple clusters. The DBP framework decomposes the optimization problem into smaller subproblems, enabling decentralized units (DUs) to solve them in parallel. To facilitate the design of the decentralized FP-based beamforming algorithm, we adopt the non-homogeneous transform and Nesterov’s extrapolation [34, 35] to avoid matrix inversion. The proposed MM-based algorithm for FA position optimization can be naturally extended to the decentralized implementation, which substantially reduces computational costs while incurring only negligible performance degradation compared to its centralized counterpart.
I-C Organization and Notation
@@
The remainder of this paper is organized as follows. Section II presents the related work. In Section III, the channel model of the FA-assisted MU-MIMO system and the formulation of the WSR maximization problem are provided. Section IV reformulates the problem using FP techniques and solves it using BCA and MM. The decentralized implementation of the proposed algorithm is introduced in Section V. Simulation results are provided in Section VI, and conclusions are drawn in Section VII.
In this paper, $a$, $\mathbf{a}$, and $\mathbf{A}$ denote a scalar, a vector, and a matrix, respectively. The imaginary unit is denoted by $\jmath$. For a complex scalar $a$, its amplitude and phase are given by $\lvert a\rvert$ and $\angle a$, respectively. The $\ell_{2}$ norm of a vector $\mathbf{a}$ is $\lVert\mathbf{a}\rVert_{2}$. $\mathbf{A}]_{m}$, $\mathbf{A}]_{mn}$, $\mathbf{A}^{\mathsf{T}}$, $\mathbf{A}^{\mathsf{H}}$, $\det(\mathbf{A})$, $\mathtt{tr}(\mathbf{A})$, $\mathrm{vec}(\mathbf{A})$, and $\lVert\mathbf{A}\rVert_{\infty}$ denote the $m$-th row, the $(m,n)$-th element, transpose, conjugate transpose, determinant, trace, vectorization, and the infinity norm of matrix $\mathbf{A}$, respectively. $\mathbf{A}\succeq\mathbf{0}$ and $\mathbf{A}\succ\mathbf{0}$ indicate that $\mathbf{A}$ is positive semi-definite and positive definite, respectively. $\mathbb{C}^{M\times N}$, $\mathbb{R}^{M\times N}$, and $\mathbb{R}_{+}^{M\times N}$ denote the sets of $M\times N$ complex, real, and non-negative real matrices, respectively. The circularly symmetric complex Gaussian (CSCG) distribution with zero mean and covariance $\sigma^{2}\mathbf{I}$ is represented as $\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I})$, and the uniform distribution over $[a,b]$ is denoted by $\mathcal{U}[a,b]$. Operator $\partial\left(\cdot\right)$ denotes the partial derivative. $\nabla_{\mathbf{x}}f\left(\mathbf{x}\right)$ and $\nabla_{\mathbf{x}}^{2}f\left(\mathbf{x}\right)$ denote the gradient vector and Hessian matrix of $f\left(\mathbf{x}\right)$ w.r.t. $\mathbf{x}$ respectively.
II Related Work
@@
The first FA position optimization algorithm was introduced in [16], where the authors considered a point-to-point MIMO system. They formulated a non-convex optimization problem and proposed an SCA-based algorithm to solve the problem. However, this work did not consider joint beamforming and FA position optimization. Due to the strong coupling between beamforming matrices and FA positions, the SCA-based approach is not directly applicable to the joint optimization problem. To address this coupling, references [23, 36] investigated joint beamforming and FA position optimization in uplink MU-MISO systems. Both works employed ZF and/or MMSE techniques to decouple beamforming and FA positions, followed by FA position optimization using multi-directional descent (MDD) and projected gradient descent (PGD), respectively. However, ZF and MMSE can lead to significant performance degradation due to their oversimplification of the objective function, and the convergence of MDD and PGD is not guaranteed. An alternative approach leverages swarm intelligence algorithms such as particle swarm optimization (PSO) [37], which offers global search capabilities and achieves better performance than MDD- and PGD-based methods. Nevertheless, PSO incurs substantially higher computational complexity, limiting its practicality for large-scale systems.
To better balance performance and complexity, reference [32] considered a multiple-input single-output single-eavesdropper (MISOSE) system and formulated secure rate maximization as a non-convex optimization problem. The authors combined the MMSE-based decoupling method with a majorization–minimization (MM) algorithm for FA position optimization. However, the overall system performance remained constrained by the MMSE method. More recently, reference [31] investigated the WSR maximization problem in an FA-assisted downlink MU-MISO system. The authors first applied the scalar WMMSE algorithm to decouple the beamforming vectors and FA positions, and then adopted the MM framework to address the resulting non-convex FA position optimization problem.
In this paper, we considered the WSR maximization problem in the more general MU-MIMO setup. Although the scalar WMMSE method [31] effectively decouples beamforming and FA position optimization with satisfactory performance in MU-MISO systems, extending this technique to MU-MIMO systems is non-trivial [38]. Instead, we adopt matrix FP techniques to handle the decoupling in the MU-MIMO setting. For FA position optimization, all surrogate functions derived in SCA- or MM-based methods [16, 32, 31] are constructed only w.r.t. a single transmit or receive FA position. These methods update one FA position at a time while keeping the others fixed, which we refer to as sequential SCA/MM. Such sequential updates result in lower computational efficiency [^3] The high computational cost of the sequential MM algorithm in [31] is due to the loop-based iterative updates. In contrast, the parallel MM algorithm proposed in this paper can be efficiently implemented using matrix operations. and is unsuitable for decentralized implementations. By contrast, the proposed parallel MM algorithm constructs a surrogate function jointly w.r.t. all transmit or receive FAs and updates them simultaneously.
III System Model and Problem Formulation
As shown in Fig. 1, we consider a downlink MU-MIMO system where a base station (BS) with $M$ FAs serves $K$ users, each equipped with $N$ FAs. A three-dimensional (3D) Cartesian coordinate system is established to describe the positions of the transmit FAs at the BS and receive FAs at the users. Specifically, let $\mathbf{t}_{m}=\left[x_{m}^{\mathtt{Tx}},y_{m}^{\mathtt{Tx}},z_{m}^{\mathtt{Tx}}\right]^{\mathsf{T}}\in\mathcal{C}_{m}^{\mathtt{Tx}}$, $1\leq m\leq M$, denote the position of $m$-th transmit FA at the BS and let $\mathbf{r}_{kn}=\left[x_{kn}^{\mathtt{Rx}},y_{kn}^{\mathtt{Rx}},z_{kn}^{\mathtt{Rx}}\right]^{\mathsf{T}}\in\mathcal{C}_{kn}^{\mathtt{Rx}}$, $1\leq k\leq K$, $1\leq n\leq N$, denote the position of the $n$-th receive FA at user $k$, where $\mathcal{C}_{m}^{\mathtt{Tx}}$ and $\mathcal{C}_{kn}^{\mathtt{Rx}}$ are the given 3D movable regions of transmit and receive FAs. Without loss of generality, the movable regions are assumed to be cuboid [23], i.e., $\mathcal{C}_{m}^{\mathtt{Tx}}=\left[x_{m}^{\min},x_{m}^{\max}\right]\times\left[y_{m}^{\min},y_{m}^{\max}\right]\times\left[z_{m}^{\min},z_{m}^{\max}\right]$ and $\mathcal{C}_{kn}^{\mathtt{Rx}}=\left[x_{kn}^{\min},x_{kn}^{\max}\right]\times\left[y_{kn}^{\min},y_{kn}^{\max}\right]\times\left[z_{kn}^{\min},z_{kn}^{\max}\right]$, $\forall k,n$, where we assume the FA movable regions of the $K$ users to be identical. In this paper, we make the following assumptions.
@@ III System Model and Problem Formulation
  • Narrow-band slow fading: The transmit and receive FAs remain static or move slowly within the movable region during each quasi-static fading block; The time required for FA movement is assumed to be much shorter than the coherence time [16].
  • Knowledge of perfect channel state information (CSI): The BS is assumed to have perfect knowledge of the downlink CSI for all users. CSI acquisition in FAS can be achieved using the methods proposed in [39, 40, 41, 42].
  • Far-field condition: The movable regions at both the users and the BS are much smaller than the signal propagation distance. The angle of departure (AoD), the angle of arrival (AoA), and the amplitude of the complex gain for each channel path are invariant to the FA positions [17].
  • Continuous FA movement: The FA positions are assumed to be continuous and can be adjusted to any location within the movable region. [^4] Due to the limited precision of stepper motors and baseband processors, motor-driven FASs involve discrete optimization at certain stages. Nevertheless, their precision is significantly higher than that of pixel-based FASs. As a result, most existing works that aim to evaluate performance limits [15, 23, 32, 31] relax the discrete optimization problem to a continuous one. We adopt the same assumption in this work.
III-A Signal Model
Let $\mathbf{s}_{k}\in\mathbb{C}^{d}$ denote the data stream intended for user $k$, where $d\leq\min\{M,N\}$ represents the number of parallel data streams. Assume that $\mathbf{s}_{k}\sim\mathcal{CN}(\mathbf{0},\mathbf{I})$. Define $\mathbf{W}_{k}\in\mathbb{C}^{M\times d}$ as the beamforming matrix for transmitting data $\mathbf{s}_{k}$ from the BS to user $k$. The received signal $\mathbf{y}_{k}$ at user $k$ is given by [^5] In this paper, we do not consider the digital combining matrix. But both the proposed centralized and decentralized algorithms can be easily extended to the case with digital combiners.